OpenStudy (anonymous):

find the derivative of y=4√x(2x+3) (square root is x only)

OpenStudy (campbell_st):

rewrite the question using indices \[y= 4x^{\frac{1}{2}}\times (2x + 3)\] you could use the product rule at this point or simplify the equation further using the index law for multiplying same base... add the powers \[y = 8x^{\frac{3}{2}} + 12x^{\frac{1}{2}}\] so you should be able to differentiate this

OpenStudy (anonymous):

ALright, I get that, however, on our exam, the answer is supposed to be 6(2x+1) all over square root of x, can you find any way to end up with this answer? Every way we've done it has given us a different answer!

OpenStudy (campbell_st):

ok... so the derivative is 12x^(1/2) + 6x^{-1/2} or \[12\sqrt{x} + \frac{6}{\sqrt{x}}\] mutliply each term by \[\sqrt{x} \] but I would have thought the simplification would have been \[\frac{6}{\sqrt{x}}(2x + 1)\] or \[\frac{6(2x +1)}{\sqrt{x}}\] and allowing for the denominator to be rationalised you would have \[\frac{6\sqrt{x}(2x + 1)}{x}\]